solve using the zero product property

(t+11)(t-5)=17

Found 2 solutions by xinxin, waynest:
Answer by xinxin(76)

(t+11)(t-5)=17
t^2 - 5t + 11t -55 =17
t^2 + 6t -72 = 0
(t+12)(t-6) = 0
Use the zero product property, t+12= 0 or t-6 =0. Therefore t = -12 or t = 6.

Answer by waynest(281)

(t +11)(t - 5) = 17
t^2 - 5t + 11t - 55 = 17
t^2 + 6t -55 -17 = 0
t^2 + 6t - 72 = 0
(t + 12)(t - 6) = 0
t + 12 = 0
t = -12
t - 6 = 0
t = 6